### v1

# RanDumb: A Simple Approach that Questions the Efficacy of Continual Representation Learning

Abstract
We propose RanDumb to examine the efficacy of continual representation learning. RanDumb embeds raw pixels using a fixed random transform which approximates an RBF-Kernel, initialized before seeing any data, and learns a simple linear classifier on top. We present a surprising and consistent finding: RanDumb significantly outperforms the continually learned representations using deep networks across numerous continual learning benchmarks, demonstrating the poor performance of representation learning in these scenarios. RanDumb stores no exemplars and performs a single pass over the data, processing one sample at a time. It complements GDumb [39], operating in a low-exemplar regime where GDumb has especially poor performance. We reach the same consistent conclusions when RanDumb is extended to scenarios with pretrained models replacing the random transform with pretrained feature extractor. Our investigation is both surprising and alarming as it questions our understanding of how to effectively design and train models that require efficient continual representation learning, and necessitates a principled reinvestigation of the widely explored problem formulation itself. Our code is available here.
1 Introduction
Continual learning is a specialized form of supervised learning, characterized by sequentially arriving tasks, coupled with additional computational and memory constraints [24, 27, 44, 49, 39] (see Verwimp et al. [51] for a survey of applications). A notable implication of this setup is the inherent performance gap between continual learning and traditional supervised learning. To elucidate this further, let us consider the achievable risk under continual learning given a dataset $\mathcal{D}$:
$\mathcal{R}_{\texttt{CL}}(\mathcal{D})\geq R_{\texttt{Joint}}(\mathcal{D})+E_{\texttt{CL}}(\mathcal{D})$ (1)
Here, $R_{\texttt{Joint}}$ denotes the risk achievable with supervised learning without any constraints, while $E_{\texttt{CL}}$ represents the gap arising from transitioning from fully supervised learning to continual learning which imposes additional constraints such as: resource-constraint (e.g., single-epoch), distribution shift, and predefined task ordering with limited access to data. Because of these additional constraints imposed by continual learning, the risk is non-decreasing and $E_{\texttt{CL}}$ is non-negative. Minimizing $E_{\texttt{CL}}$ has been the primary focus of the continual learning literature [16, 29, 37, 41].
Figure 1: Mechanism of RanDumb. Project raw pixels to a high dimensional space using random fourier projections ($\varphi$), then decorrelate the features using Mahalanobis distance [31] and classify with the nearest class mean. The online update only involves updating a single covariance matrix and class-means.
In this work, we discover that the fundamental aspect of continual representation learning (CRL) is often overlooked in the continual learning literature, as most methods merge continual classifier and representation learning. We dissect these two aspects and study the effect of representation learning on $E_{\texttt{CL}}$. The central question of our work is trying to answer:
Are current continual learning setups overly constrained for effective continual representation learning?
Specifically, consider a continual representation learning algorithm $\mathtt{CRL}$ which learns a representation from the continual dataset $\mathcal{D}$. Is this representation better than a random (ie. non-learned) representation RandRep? More formally, is the following quantity positive or negative?
$RepGain(\mathcal{D})=E_{\texttt{CL}}(\mathcal{D},\mathtt{RandRep})-E_{\texttt{CL}}(\mathcal{D},\mathtt{CRL}(\mathcal{D}))$
Here, the random representation functions are defined at initialization without seeing any data or using domain-specific priors. Note that one expects the impact of continual representation learning ($RepGain(\mathcal{D})$) to be positive, as continual representation learning approaches should improve performance over even the best random representation.
Our primary contribution is empirically demonstrating that currently $RepGain(\mathcal{D})<0$, i.e. there exists random representations which outperform state-of-the-art representations learned by continual learning algorithms using deep networks. We show this by introducing a straightforward baseline, which we name RanDumb, standing for Random representation function and a Dumb linear classifier. Despite replacing the deep embedder with a random function, we see from Table 1 (top) that RanDumb outperforms the current state-of-the-art methods being exemplar-free over challenging large task scenarios given in [59]. We extend the large task scenario to methods with pretrained feature extractors [64]. In RanDumb, we replace our random projection with the pretrained feature extractor here, simply learning a linear classifier. We show results in Table 2 (top), showing that RanDumb outperforms the best methods in this scenario as well.
Table 1: Online Continual Learning. Performance comparison of RanDumb on the PEC setup [59] and VAE-GC [50]. Setup and numbers borrowed from PEC [59]. RanDumb outperforms the best exemplar-free OCL method. Comparing representation learning with no representation learning, a fixed, random function outperforms training a deep network continually, i.e. $RepGain(\mathcal{D})<0$. Furthermore, RanDumb nearly matches performance of one-pass joint, demonstrating the inefficacy of current benchmarks.
Method MNIST CIFAR10 CIFAR100 m-ImNet
Comparison with Best Method
Best (PEC) 92.3 58.9 26.5 14.9
RanDumb (Ours) 98.3 55.6 28.6 17.7
Improvement +6.0 -3.3 +2.1 +2.8
Only Ablating Representation Learning: Random v/s Deep Embedding
VAE-GC 84.0 42.7 19.7 12.1
RanDumb (Ours) 98.3 55.6 28.6 17.7
Error ($E_{\texttt{CRL}}$) +14.3 +12.9 +8.9 +5.6
Scope of Improvement
Joint (One Pass) 98.3 74.2 33.0 25.3
RanDumb (Ours) 98.3 55.6 28.6 17.7
Gap Covered. (%) 100% 75% 87% 70%
Table 2: Offline Continual Learning. Performance comparison with ImageNet21K ViT-B16 model using 2 init classes and 1 new class per task. RanDumb here assumes the features as input pixels. SLCA [62] and RanPAC [30] results reproduced from original codebase. Comparing representation learning with no representation learning, a fixed, pretrained feature extractor outperforms further finetuning a deep network continually, i.e. $RepGain(\mathcal{D})<0$. RanDumb nearly matches performance of joint, demonstrating the inefficacy of current benchmarks. *here representation learning collapses, accuracy reported without PETL. ${}^{+}$here representation learning and random projection both collapse, reporting last stable accuracy.
Method CIFAR IN-A IN-R CUB OB VTAB Cars
Comparison with Best Method
Best (RanPAC) 89.6 ${}^{+}$ 67.3 ${}^{+}$ 77.9* ${}^{+}$ $*{}^{*}$
RanDumb (Ours) 86.8 42.2 64.9 88.5 75.3 92.3 69.1
Improvement -2.9 +15.4 -2.4 +1.3 -2.6 +3.7 +15.4
Only Ablating Representation Learning: No Finetuning v/s Full Finetuning
SLCA 86.8 - 54.2 82.1 - - 18.2
RanDumb (Ours) 86.8 42.2 64.9 88.5 75.3 92.3 69.1
Improvement +0.0 - +10.7 +6.4 - - +50.9
Scope of Improvement
Joint 93.8 70.8 86.6 91.1 83.8 95.5 86.9
RanDumb (Ours) 86.8 42.2 64.9 88.5 75.3 92.3 69.1
Gap Covered. (%) 93% 60% 75% 97% 92% 97% 80%
Figure 2: Mechanism of RanDumb. RanDumb projects the input datapoints to a high-dimensional space to create a clearer separation between classes. Subsequently, it corrects the anisotropy across different feature dimensions, scaling them to be unit variance each. This allows a nearest class mean classifier to accurately rely on cosine similarity. The figure is adapted from Pilario et al. [38].
1.1 RanDumb: Construction and Performance
Mechanism. Our mechanism is illustrated in Figure 1. RanDumb first projects input pixels into a high-dimensional space using a fixed kernel based on random Fourier basis, which is a low-rank data-independent approximation of the RBF Kernel [42]. Then, we use a simple linear classifier which first normalizes distances across different feature dimensions (anisotropy) with Mahalanobis distance [31] and then uses nearest class means for classification [33]. In scenarios with pretrained feature extractors, we use the fixed pretrained model as embedder and learn a linear classifier over it, similar to Hayes and Kanan [20].
Properties. RanDumb needs no storage of exemplars and requires only one pass over the data in a one-sample-per-timestep fashion. Furthermore, it only requires online estimation of the covariance matrix and nearest class mean, hence is invariant to data order, representing the accuracy in the worst-case ordering.
Poor Representation Learning. We first compare RanDumb with VAE-GC [50] in Table 1 (middle) and SLCA [62] in Table 2 (middle). The only distinction between them is their embedding functions: RanDumb uses a fixed function (random/pretrained network) for the embedder, whereas VAE-GC and SLCA further continually trained deep networks, using them as the embedder. Astonishingly, RanDumb consistently surpasses VAE-GC and SLCA baselines by wide margins of 5-15%. This shows that state-of-the-art online continual learning algorithms fail to learn effective representations in a single pass across standard online continual learning datasets.
Benchmarks Over-Constrained. The reader might now ask: With better representation learning, can one outperform RanDumb by large margins in these settings? To study this, we compare the performance of RanDumb with Joint in both online and offline settings respectively in Table 1 (bottom) and Table 2 (bottom). We observe that our simple baseline RanDumb recovers 70-90% of the gap to the respective joint classifiers across both settings. Improving forgetting in continual representation learning [21, 36] offers limited scope for improvement, as current benchmarks might be too constrained to have effective representation learning at all.
In the next sections, we shall first describe our proposed method and then walk through our findings in detail.
2 RanDumb: Mechanism & Intuitions
In this section, we thoroughly examine RanDumb, focusing on its two main elements: random projection and the dumb learner. We illustrate the mechanism of RanDumb using three toy examples in Figure 2.
We first describe the dumb classifier. To classify a test sample $\bf x_{\rm test}$, we start with a simple classifier, the nearest class mean (NCM) classifier. It predicts the class among $C$ classes by highest value of the similarity function $f$ among class means $\mathbf{\mu}_{i}$:
$y_{\textrm{pred}}=\operatorname*{arg\,max}_{{i}\in\{1,\ldots,|C|\}}f({\bf x_{\textrm{test}}},\mathbf{\mu}_{i}),\text{where}$
$f({\bf x_{\textrm{test}}},\mathbf{\mu}_{i}):={\bf x_{\textrm{test}}}^{\top}{\mathbf{\mu}_{i}}$ (2)
and $\mathbf{\mu}_{i}$ are the class-means in the pixel space:
$\mathbf{\mu}_{i}=\frac{1}{|C_{i}|}\sum_{\mathbf{x}}\mathbf{x}\qquad\forall{\bf x}\in C_{i}$
RanDumb adds two additional components to this classifier: 1) Kernelization and 2) Decorrelation.
1
Kernelization: Why & How? Classes are typically not linearly separable in the pixel space, unlike in the feature space of deep models. Hence, we apply the kernel trick to embed the pixels in a better representation space, computing all distances between the data and class-means in this embedding space. This phenomena is illustrated on three toy examples to build intuitions in Figure 2 (Embed). We use an RBF-Kernel, which for two points $\mathbf{x}$ and $\mathbf{y}$ is defined as: $K_{\text{RBF}}(\mathbf{x},\mathbf{y})=\exp\left(-\gamma\|\mathbf{x}-\mathbf{y}\|^{2}\right)$ where $\gamma$ is a scaling parameter. However, calculating the RBF kernel is not possible due to the online continual learning constraints preventing computation of pairwise-distance between all points. Hence, we use a data-independent approximation, random Fourier projection $\phi(\mathbf{x})$, as given in [42]:
$K_{\text{RBF}}(\mathbf{x},\mathbf{y})\approx\phi(\mathbf{x})^{T}\phi(\mathbf{y})$
[…]
$\phi(\mathbf{x})=\frac{1}{\sqrt{D}}\left[\cos(\mathbf{\omega}_{1}^{T}\mathbf{x}),\sin(\mathbf{\omega}_{1}^{T}\mathbf{x}),..,\cos(\mathbf{\omega}_{D}^{T}\mathbf{x}),\sin(\mathbf{\omega}_{D}^{T}\mathbf{x})\right]$
We keep these $\omega$ bases fixed throughout online learning. Thus, we obtain our modified similarity function from Equation 2 as:
$f({\bf x_{\rm test}},\mathbf{\mu}_{i}):={\bf\phi(x_{\rm test})}^{\top}{\bar{\mathbf{\mu}}_{i}}$ (3)
where $\bar{\mathbf{\mu}}_{i}$ are the class-means in the kernel space:
$\bar{\mathbf{\mu}}_{i}=\frac{1}{|C_{i}|}\sum_{\mathbf{x}}\phi(\mathbf{x})\qquad\forall{\bf x}\in C_{i}$
2
Decorrelation: Why & How? Projected raw pixels, similar to features of deep models, have feature dimensions with different variances (anisotropic). Hence, instead of naively computing ${\bf\phi(x_{\rm test})}^{\top}{\bar{\mathbf{\mu}}_{i}}$, we further decorrelate the feature dimensions using a Mahalonobis distance with the empirical covariance matrix $\mathbf{S}$. We illustrate this phenomena as well on three toy examples in Figure 2 (Decorrelate) to build intuitions. Our similarity function finally is:
$f({\bf x_{\rm test}},\mathbf{\mu}_{i}):=(\phi(\mathbf{x}_{\rm test})-\bar{\mathbf{\mu}}_{i})^{T}\mathbf{S}^{-1}(\phi(\mathbf{x}_{\rm test})-\bar{\mathbf{\mu}}_{i})$ (4)
Online Computation. Our random projection is computed and fixed before seeing any real data. In the continual learning process, we only update the running class mean and empirical covariance matrix, which is accurately estimable online [^1] Online prediction is possible using the Sherman–Morrison formula to update the inverse of the covariance matrix. However, we only needed to compute inverse at the last timestep for all our experiments, hence we do not experiment with this..
Note on Equivalences. For the curious reader we draw equivalences between the linear classifier used in this work and related works. These equivalences rely on the assumption that the classes are equiprobable, which is the case for most datasets here. In these cases, nearest class mean classifier with the Mahalanobis distance metric is equivalent to linear discriminant analysis classifier [32]. Hence, one could say RanDumb is equivalent to a Streaming LDA classifier with an approximate RBF Kernel. Alternatively, one could think of the decorrelation operation as explicitly decorrelating the features with ZCA whitening [7]. We hope these alternative perspectives allow for a better understanding of the RanDumb classifier.
Table 3: Benchmark (Overview). We evaluate RanDumb across a diverse set of benchmarks that progressively relax constraints compared to RanDumb. Benchmark A closely matches RanDumb with one class per timestep and no stored exemplars. Benchmark B, D, E progressively relax the constraints on exemplars and classes per timestep. Benchmark C and E remove the online constraint by allowing unrestricted training and sample access within a task without exemplar-storage of past tasks. Benchmark F allows using large pretrained models, modified by us with one class per task, inspired by challenges detailed in Benchmark A.
Setup Num #Classes #Samples #Stored Contrastive
Passes Per Task Per Step Exemplars Augment
RanDumb (Ours) 1 1 1 0 No
A (Zając et al. [59]) 1 1 10 0 No
B1 (Guo et al. [18]) 1 2 10 100-2000 No
B2 (Guo et al. [18]) 1 2 10 100-1000 Yes
C (Smith et al. [48]) Many 10 All 0 No
D (Anonymous [4]) 1 2-10 10 1000 No
E (Ye and Bors [58]) 1 2-5 10 1000-5000 No
F (Wang et al. [54]) Many 1 All 0 No
Table 4: Benchmark A (Ref: Table 1 from PEC [59]). We perform comprehensive comparisons of RanDumb with popular online continual learning approaches in a 1-class per task setting referred as ‘Dataset (num_tasks/1)’. This setting is a challenging setting for OCL methods except GDumb. We observe that RanDumb outperforms all approaches across all datasets by 2-6% margins, with an exception of latest work PEC [59] on CIFAR10.
Method Memory MNIST (10/1) CIFAR-10 (10/1) CIFAR-100 (100/1) miniImageNet (100/1)
Fine-tuning all 10.1$\pm$ 0.0 10.0$\pm$ 0.0 1.0$\pm$ 0.0 1.0$\pm$ 0.0
Joint, 1 epoch all 98.3$\pm$ 0.0 74.2$\pm$ 0.1 33.0$\pm$ 0.2 25.3$\pm$ 0.2
Rehearsal-Based
ER [13] (ICML-W ’19) 500 84.4$\pm$ 0.3 40.6$\pm$ 1.1 12.5$\pm$ 0.3 5.7$\pm$ 0.2
A-GEM [12] (ICLR ’19) 500 59.8$\pm$ 0.8 10.2$\pm$ 0.1 1.0$\pm$ 0.0 1.1$\pm$ 0.1
iCaRL [44] (CVPR ’17) 500 83.1$\pm$ 0.3 37.8$\pm$ 0.4 5.7$\pm$ 0.1 7.5$\pm$ 0.1
BiC [56] (CVPR ’19) 500 86.0$\pm$ 0.4 35.9$\pm$ 0.4 6.4$\pm$ 0.3 1.5$\pm$ 0.1
ER-ACE [10] (ICLR ’22) 500 87.8$\pm$0.2 39.9$\pm$0.5 8.2$\pm$0.2 5.7$\pm$0.2
DER [9] (NeurIPS ’20) 500 91.7$\pm$ 0.1 40.0$\pm$ 1.5 1.0$\pm$ 0.1 1.0$\pm$ 0.0
DER++ [9] (NeurIPS ’20) 500 91.9$\pm$ 0.2 35.6$\pm$ 2.4 6.2$\pm$ 0.4 1.4$\pm$ 0.1
X-DER [8] (TPAMI ’22) 500 83.0$\pm$ 0.1 43.2$\pm$ 0.5 15.6$\pm$ 0.1 8.2$\pm$ 0.4
GDumb [39] (ECCV ’20) 500 91.0$\pm$0.2 50.7$\pm$0.7 8.2$\pm$0.2 -
Rehearsal-Free
EWC [24] (PNAS ’17) 0 10.1$\pm$ 0.0 10.6$\pm$ 0.4 1.0$\pm$ 0.0 1.0$\pm$ 0.0
SI [60] (ICML ’17) 0 12.7$\pm$ 1.0 10.1$\pm$ 0.1 1.1$\pm$0.0 1.0$\pm$0.1
LwF [26] (TPAMI ’17) 0 11.8 $\pm$ 0.6 10.1$\pm$ 0.1 0.9$\pm$0.0 1.0$\pm$ 0.0
LT [61] (Arxiv ’18) 0 10.9$\pm$ 0.9 10.0$\pm$ 0.2 1.1$\pm$ 0.1 1.0$\pm$ 0.0
Gen-NCM [22] (NeurIPS-W ’22) 0 82.0$\pm$ 0.0 27.7$\pm$ 0.0 10.0$\pm$ 0.0 7.5$\pm$ 0.0
Gen-SLDA [20] (CVPR-W ’20) 0 88.0$\pm$ 0.0 41.4$\pm$ 0.0 18.8$\pm$ 0.0 12.9$\pm$ 0.0
VAE-GC [50] (CVPR-W ’21) 0 84.0$\pm$ 0.5 42.7$\pm$ 1.3 19.7$\pm$ 0.1 12.1$\pm$ 0.1
PEC [59] (ICLR ’24) 0 92.3$\pm$ 0.1 58.9$\pm$ 0.1 26.5$\pm$ 0.1 14.9$\pm$ 0.1
RanDumb (Ours) 0 98.3 (+5.9) 55.6 (-3.3) 28.6 (+2.1) 17.7 (+2.8)
3 Experiments
We compare RanDumb with algorithms across online continual learning benchmarks with an emphasis on exemplar-free and low-exemplar storage regime complementing GDumb [39] in regimes where it has poor performance.
Benchmarks. We illustrate the benchmarks used in our experiments along with key differences to training run for RanDumb in Table 3. We aim for a truly comprehensive coverage and show results on four different benchmarks (A, B, D, E) which reflect the latest trends in online continual learning (’22-’24) across exemplar-free, contrastive-training [^2] Benchmark B is split into two distinct sections: (i) methods that do not rely on contrastive learning and heavy augmentation, named B1, and (ii) approaches that incorporate contrastive learning and extra augmentations, named as Scenario B2., meta-continual learning and network-expansion based approaches respectively and a rehearsal-free offline continual learning benchmark C. These benchmarks are ordered by increasingly relaxed constraints, moving further away from the training scenario of RanDumb. These benchmarks train models from scratch. Hence, RanDumb utilizes a random Fourier transform for embedding.
We further test on exemplar-free scenarios in offline continual learning using Benchmark F [54] with the challenging one-class per task constraint borrowed from [59]. This benchmark allows using pretrained models along with unrestricted training time and access to all class samples at each timestep. However, RanDumb is restricted to learning from a single pass seeing only one sample at a time. RanDumb only learns a linear classifier over a given pretrained model in Benchmark F. We use LAMDA-PILOT codebase [^3] https://github.com/sun-hailong/LAMDA-PILOT for all methods, except RanPAC and SLDA which use their codebase. All methods use the original hyperparameters. We only change initial classes to 2 and number of classes per task to 1 in the scripts and test using both ImageNet21K and ImageNet1K ViT-B/16 models. For detailed assumptions, task orderings, implementation details and other specific information of these benchmarks and approaches, readers are encouraged to consult the LAMDA-PILOT codebase.
Implementation Details (RanDumb). We evaluate RanDumb using five datasets commonly used for online continual learning: MNIST, CIFAR10, CIFAR100, TinyImageNet200, and miniImageNet100. For the latter two datasets, we downscale all images to 32x32. We augment each datapoint with flipped version, hence two images are seen by the classifier at each timestep (except for MNIST and Benchmark F). We normalize all images and flatten them into vectors, obtaining 784-dim input vectors for MNIST and 3072-dim input vectors for all the other datasets. For Benchmark F, we compare RanDumb on seven datasets used in LAMDA-PILOT, replacing ObjectNet with Stanford Cars as ObjectNet license prohibits training models on it. We use the 768-dimensional features from the same pretrained ViT-B models used in this benchmark. We measure accuracy on the test set of all past seen classes after completing the full one-pass over the dataset. We take the average accuracy after the last task on all past tasks [59, 18, 54]. In Benchmark A and F, since we have one class per task, the average accuracy across past tasks is the same regardless of the task ordering. In Benchmarks A-E, all datasets have the same number of samples, hence similarly the average accuracy across past tasks is the same regardless of the task ordering. We used the Scikit-Learn implementation of Random Fourier Features [42] with 25K embedding size, $\gamma=1.0$. We use progressively increasing ridge regression parameter ($\lambda$) with dataset complexity, using $\lambda=10^{-6}$ for MNIST, $\lambda=10^{-5}$ for CIFAR10/100 and $\lambda=10^{-4}$ for TinyImageNet200/miniImageNet100. All experiments were conducted on a CPU server with a 48-core Intel Xeon Platinum 8268 CPU and 392GB of RAM. Our code is available at https://github.com/drimpossible/RanDumb.
Table 5: Benchmark B.1 (Ref: Table adopted from OnPro [55], OCM[18]) We perform comprehensive comparisons of RanDumb with popular online continual learning approaches in less ideal many-classes per task setting referred as ‘Dataset (num_tasks/num_classes_per_task)’. Furthermore, results are categorized by memory buffer sizes given in ‘M’ subcolumn for that dataset. Comparing across rows, we observe that RanDumb outperforms the best among the compared approaches without heavy-augmentations by 3-20% margins despite being exemplar free. Only in one case, it is second best, but after GDumb. Comparing the two columns of CIFAR100, we observe a large drop in performance of OCL methods on benchmarks with longer timesteps indicating analyzing longer timesteps is important in continual learning.
Method MNIST (5/2) CIFAR10 (5/2) CIFAR100 (10/10) CIFAR100 (50/2) TinyImageNet (100/2)
$M=0.1k$ $M=0.1k$ $M=0.2k$ $M=0.5k$ $M=1k$ $M=1k$ $M=1k$ $M=2k$
AGEM [12] (ICLR ’19) 56.9±5.2 17.7±0.3 22.7±1.8 5.8±0.2 5.9±0.1 1.8±0.2 0.8±0.1 0.9±0.1
GSS [3] (NeurIPS ’19) 70.4±1.5 18.4±0.2 26.9±1.2 8.1±0.2 11.1±0.2 4.3±0.2 1.1±0.1 3.3±0.5
ER [13] (ICML-W ’19) 78.7±0.4 19.4±0.6 29.7±1.0 8.7±0.3 15.7±0.3 8.3±0.3 1.2±0.1 5.6±0.5
ASER [46] (AAAI ’21) 61.6±2.1 20.0±1.0 27.8±1.0 11.0±0.3 16.4±0.3 9.6±1.3 2.2±0.1 5.3±0.3
MIR [2] (NeurIPS ’19) 79.0±0.5 20.7±0.7 37.3±0.3 9.7±0.3 15.7±0.2 12.7±0.3 1.4±0.1 6.1±0.5
ER-AML [10] (ICLR ’22) 76.5±0.1 - 40.5±0.7 - 16.1±0.4 - - 5.4±0.2
iCaRL [44] (CVPR ’17) - 31.0±1.2 33.9±0.9 12.8±0.4 16.5±0.4 - 5.0±0.3 6.6±0.4
DER++ [9] (NeurIPS ’20) 74.4±1.1 31.5±2.9 44.2±1.1 16.0±0.6 21.4±0.9 9.3±0.3 3.7±0.4 5.1±0.8
GDumb [39] (ECCV ’20) 81.2±0.5 23.3±1.3 35.9±1.1 8.2±0.2 18.1±0.3 18.1±0.3 4.6±0.3 12.6±0.1
CoPE [15] (CVPR ’21) - 33.5±3.2 37.3±2.2 11.6±0.4 14.6±1.3 - 2.1±0.3 2.3±0.4
DVC [18] (CVPR ’22) - 35.2±1.7 41.6±2.7 15.4±0.3 20.3±1.0 - 4.9±0.6 7.5±0.5
Co²L [11] (ICCV ’21) 83.1±0.1 - 42.1±1.2 - 17.1±0.4 - - 10.1±0.2
R-RT [6] (CVPR ’21) 89.1±0.3 - 45.2±0.4 - 15.4±0.3 - - 6.6±0.3
CCIL [35] (CVPR ’21) 86.4±0.1 - 50.5±0.2 - 18.5±0.3 - - 5.6±0.9
IL2A [65] (NeurIPS ’21) 90.2±0.1 - 54.7±0.5 - 18.2±1.2 - - 5.5±0.7
BiC [56] (CVPR ’19) 90.4±0.1 - 48.2±0.7 - 21.2±0.3 - - 10.2±0.9
SSIL [1] (ICCV ’21) 88.2±0.1 - 49.5±0.2 - 26.0±0.1 - - 9.6±0.7
Rehearsal-Free
PASS [65] (NeurIPS ’21) - 33.7±2.2 33.7±2.2 7.5±0.7 7.5±0.7 - 0.5±0.1 0.5±0.1
RanDumb (Ours) 98.3 (+7.8) 55.6 (+20.4) 55.6 (+5.9) 28.6 (+12.6) 28.6 (+2.6) 28.6 (+10.5) 11.6 (+6.6) 11.6 (-1.0)
Table 6: (Left) Benchmark B.2 (Ref: Table from OnPro [55]) We now compare with state-of-the-art contrastive representation learning based online continual learning approaches which additionally use sophisticated augmentations. These augmentations and additional loss function can improve any of the OCL methods described in Setup B.1 by large margins [18]. We observe that RanDumb often outperforms these sophisticated methods despite all of these factors on small-exemplar settings. (Right) Benchmark C (Ref: Table 2 from [48]). We compare the performance of RanDumb with latest efforts on improving rehearsal-free methods. We outperform them by 4% margins.
Method MNIST (5/2) CIFAR10 (5/2) CIFAR100 (10/10) TinyImageNet (100/2)
$M=0.1k$ $M=0.1k$ $M=0.5k$ $M=1k$
SCR [28] (CVPR ’21) 86.2±0.5 40.2±1.3 19.3±0.6 8.9±0.3
OCM [18] (ICML ’22) 90.7±0.1 47.5±1.7 19.7±0.5 10.8±0.4
OnPro [55] (ICCV ’23) - 57.8±1.1 22.7±0.7 11.9±0.3
Rehearsal-Free
RanDumb (Ours) 98.3 (+7.5) 55.6 (-2.2) 28.6 (+5.9) 11.6 (-0.3)
Method CIFAR100
(10/10)
Rehearsal-Free
PredKD [26] 24.6
PredKD + FeatKD 12.4
PredKD + EWC 23.3
PredKD + L2 21.5
RanDumb 28.6 (+4.0)
Table 7: (Left) Benchmark D (Ref: Table 2 from VR-MCL [4]) We compare RanDumb with meta-continual learning approaches operating in a high memory setting, allowing buffer sizes up to 1K exemplars. Despite catering representations specifically to combat forgetting, all methods with the exception of VR-MCL are outperformed by RanDumb on TinyImageNet. RanDumb also surpasses all prior work by a substantial 9.1% on CIFAR100. Allowing generous replay buffers shifts scenarios to a high exemplar regime where GDumb performs the best on CIFAR10. Yet RanDumb competes favorably even under these conditions not optimized for its approach. (Middle) Benchmark E (Ref: Table 1 from SEDEM [58]) We compare RanDumb with recent network expansion based online continual learning approaches. Despite allowing access to much larger memory buffers, RanDumb matches the performance of best method SEDEM on MNIST, while exceeding it by 0.3% on CIFAR10 and 3.8% on CIFAR100. (Right) Architectures (Ref: Table 1 from Mirzadeh et al. [34]) RanDumb surpasses continual representation learning across a wide range of architectures, achieving close to 94% of the joint performance.
Method CIFAR10 CIFAR100 TinyImageNet
(5/2) (10/10) (20/10)
$M=1k$ $M=1k$ $M=1k$
Finetune 17.0 $\pm$ 0.6 5.3 $\pm$ 0.3 3.9 $\pm$ 0.2
A-GEM [12] (ICLR ’19) 18.4 $\pm$ 0.2 6.0 $\pm$ 0.2 4.0 $\pm$ 0.2
IS [60] (ICML ’17) 17.4 $\pm$ 0.2 5.2 $\pm$ 0.2 3.3 $\pm$ 0.3
MER [45] (ICLR ’19) 36.9 $\pm$ 2.4
La-MAML [19] (NeurIPS ’20) 33.4 $\pm$ 1.2 11.8 $\pm$ 0.6 6.74 $\pm$ 0.4
GDumb [39] (ECCV ’20) 61.2 $\pm$ 1.0 18.1 $\pm$ 0.3 4.6 $\pm$ 0.3
ER [13] (ICML-W ’19) 43.8 $\pm$ 4.8 16.1 $\pm$ 0.9 11.1 $\pm$ 0.4
DER [9] (NeurIPS ’20) 29.9 $\pm$ 2.9 6.1 $\pm$ 0.1 4.1 $\pm$ 0.1
DER++ [9] (NeurIPS ’20) 52.3 $\pm$ 1.9 11.8 $\pm$ 0.7 8.3 $\pm$ 0.3
CLSER [5] (ICLR ’22) 52.8 $\pm$ 1.7 17.9 $\pm$ 0.7 11.1 $\pm$ 0.2
OCM [18] (ICML ’22) 53.4 $\pm$ 1.0 14.4 $\pm$ 0.8 4.5 $\pm$ 0.5
ER-OBC [14] (ICLR ’23) 54.8 $\pm$ 2.2 17.2 $\pm$ 0.9 11.5 $\pm$ 0.2
VR-MCL [4] (ICLR ’24) 56.5 $\pm$ 1.8 19.5 $\pm$ 0.7 13.3 $\pm$ 0.4
Rehearsal-Free
RanDumb (Ours) 55.6 (-5.6) 28.6 (+9.1) 11.6 (-1.7)
Method MNIST CIFAR10 CIFAR100
(5/2) (5/2) (20/5)
$M=2k$ $M=1k$ $M=5k$
Finetune 19.8 $\pm$ 0.1 18.5 $\pm$ 0.3 3.5 $\pm$ 0.1
MIR [2] (NeurIPS ’19) 93.2 $\pm$ 0.4 42.8 $\pm$ 2.2 20.0 $\pm$ 0.6
GEM [12] (ICLR ’19) 93.2 $\pm$ 0.4 24.1 $\pm$ 2.5 11.1 $\pm$ 2.4
iCARL [44] (CVPR ’17) 83.9 $\pm$ 0.2 37.3 $\pm$ 2.7 10.8 $\pm$ 0.4
G-MED [23] (NeurIPS ’21) 82.2 $\pm$ 2.9 47.5 $\pm$ 3.2 19.6 $\pm$ 1.5
GSS [3] (NeurIPS ’19) 92.5 $\pm$ 0.9 38.5 $\pm$ 1.4 13.1 $\pm$ 0.9
CoPE [15] (CVPR ’21) 93.9 $\pm$ 0.2 48.9 $\pm$ 1.3 21.6 $\pm$ 0.7
CURL [43] (NeurIPS ’19) 92.6 $\pm$ 0.7 - -
CNDPM [25] (ICLR ’20) 95.4 $\pm$ 0.2 48.8 $\pm$ 0.3 22.5 $\pm$ 1.3
Dynamic-OCM [57] (ECCV ’22) 94.0 $\pm$ 0.2 49.2 $\pm$ 1.5 21.8 $\pm$ 0.7
SEDEM [58] (ICCV ’23) 98.3 $\pm$ 0.2 55.3 $\pm$ 1.3 24.8 $\pm$ 1.2
Rehearsal-Free
RanDumb (Ours) 98.3 (0.0) 55.6 (+0.3) 28.6 (+3.8)
Model CIFAR100
Joint 79.58
CNN x1 62.2 $\pm$1.35
CNN x2 66.3 $\pm$1.12
CNN x4 68.1 $\pm$0.5
CNN x8 69.9 $\pm$0.62
CNN x16 76.8 $\pm$0.76
ResNet-18 45.0 $\pm$0.63
ResNet-34 44.8 $\pm$2.34
ResNet-50 56.2 $\pm$0.88
ResNet-101 56.8 $\pm$1.62
WRN-10-2 50.5 $\pm$2.65
WRN-10-10 56.8 $\pm$2.03
WRN-16-2 44.6 $\pm$2.81
WRN-16-10 51.3 $\pm$1.47
WRN-28-2 46.6 $\pm$2.27
WRN-28-10 49.3 $\pm$2.02
ViT-512/1024 51.7 $\pm$1.4
ViT-1024/1546 60.4 $\pm$1.56
RandDumb (Ours) 74.8 (-2.0)
3.1 Results
We extensively evaluate RanDumb and detail our findings on each benchmark in this section.
Benchmark A. Benchmark A assesses continual learning models in the challenging setup of one class per timestep, closely mirroring our training assumptions. We present our results in Table 4. Comparing across rows, and see that RanDumb improves over prior state-of-the-art across all datasets with 2-6% margins. The only exception is PEC on CIFAR10, where RanDumb underperforms by 3.3%. Nonetheless, it still outperforms the second-best model, GDumb with a 500 memory size, by 4.9%.
Benchmark B1. We present our results comparing with non-contrastive methods in Table 5. We notice that scenario allows two classes per task and relaxes the memory constraints for online continual learning methods, allowing for higher accuracies compared to Benchmark A. Despite that, RanDumb outperforms latest OCL algorithms on MNIST, CIFAR10 and CIFAR100—often by margins exceeding $>$10%. The lone exception is GDumb achieving a higher performance with 2K memory samples on TinyImageNet, indicating that this already is in the high-memory regime. Overall, RanDumb is the best algorithm across the benchmark.
Benchmark B2. We additionally compare our performance with the latest online CL approaches using contrastive losses with sophisticated data augmentations. As shown in in Table 6 (Left), these advancements provide large performance improvements over methods from Benchmark B.1. To compensate, we compare on lower exemplar budgets. The best approach, OnPro [55], which outpeforms RanDumb on CIFAR10 by 2.2% and TinyImageNet by 0.3%, but falls significantly short on CIFAR100 by 5.9%. Overall, RanDumb achieves strong results compared to highly performant representation learning using state-of-the-art contrastive learning approaches customized to continual learning, despite storing no exemplars.
Benchmark C. We now compare against offline rehearsal-free continual learning approaches in Table 6 (Right) on CIFAR100 dataset. Despite online training, RanDumb outperforms PredKD by over 4% margins.
Benchmark D. We now compare performance of RanDumb against meta-continual learning methods, which require large exemplars with buffer sizes of 1K in Table 7 (left). RanDumb achieves strong performance under these conditions, exceeding all prior work by a large margin of 9.1% on CIFAR100 and outperforms all but VR-MCL approach on the TinyImageNet dataset. GDumb performs the best on CIFAR10, indicating this is already in a large-exemplar regime uniquely unsuited for RanDumb.
Benchmark E. We compare RanDumb against recent network expansion-based online continual learning methods in Table 7 (middle). These approaches grow model capacity to mitigate forgetting while dealing with shifts in the data distribution, and are allow larger memory buffers. RanDumb matches the performance of the state-of-the-art method SEDEM [58] on MNIST, while exceeding it by 0.3% on CIFAR10 and 3.8% on CIFAR100. The results indicate current benchmarks remain too restrictive for effective continual representation learning.
Benchmark F. We compare performance of approaches which do not further train the deep network like RanDumb (ours) [^4] Recall that RanDumb in Benchmark F is quite similar to SLDA [20]. and NCM [22] against popular prompt-tuning approaches in Table 8. and discover that prompt-tuning approaches completely collapse under large timesteps and approaches which do not finetune their pretrained model achieve strong performance, even under challenging one class per timestep constraint. Note that there are variants of RanDumb (e.g. RanPAC - RP - PETL [30]) that achieve higher accuracies due to minor modifications which are designed for this benchmark.
Conclusion. Overall, despite RanDumb being exemplar-free and modeling worst-case ordering, it outperforms nearly all online continual learning methods across various tasks when exemplar storage is limited. We specifically benchmark on lower exemplar sizes to complement settings in which GDumb does not perform well.
Table 8: Benchmark F We compare RanDumb with other prompt-tuning based continual learning approaches using ViT-B/16 ImageNet-21K/1K pretrained models using 2 init classes and 1 class per task setting. Most prompt-tuning based methods collapse and simple baselines which do not pretrain the architecture like NCM [22] or RanDumb achieve state-of-the-art performance. ’-’ indicates that despite reasonable efforts, we could not run the method. ${}^{+}$here representation learning and random projection both collapse, reporting last stable accuracy.
Method CIFAR IN-A IN-R CUB VTAB
ViT-B/16 (IN-1K Pretrained)
Finetune 1.0 1.2 1.1 1.0 2.1
L2P [54] (CVPR ’22) 2.4 0.3 0.8 1.4 1.3
DualPrompt [53] (ECCV ’22) 2.3 0.3 0.8 0.9 4.2
CODA-Prompt [47] (CVPR ’23) 2.6 0.3 0.8 1.9 6.3
Adam-Adapt [64] (Arxiv ’23) 76.7 49.3 62.0 85.2 83.6
Adam-SSF [64] (Arxiv ’23) 76.0 47.3 64.2 85.6 84.2
Adam-VPT [64] (Arxiv ’23) 79.3 35.8 61.2 83.8 86.9
Adam-FT [64] (Arxiv ’23) 72.6 49.3 61.0 85.2 83.8
Memo [63] (ICLR ’23) 69.8 - - 81.4 -
iCARL [44] (CVPR ’17) 72.4 - 35.2 72.4 -
Foster [52] (ECCV ’22) 52.2 - 76.8 86.6 -
NCM [22] (NeurIPS-W ’22) 76.2 49.4 61.2 85.2 83.6
SLCA [62] (ICCV ’23) 86.3 - 52.8 84.7 -
RanPAC [30] (NeurIPS ’23) 88.2 ${}^{+}$ 72.8 ${}^{+}$ 93.0
RanDumb (Ours) 84.5 47.8 66.9 88.0 93.4
ViT-B/16 (IN-21K Pretrained)
Finetune 2.8 0.5 1.2 1.2 0.5
Adam-Adapt [64] (Arxiv ’23) 82.4 48.8 55.4 86.7 84.4
Adam-SSF [64] (Arxiv ’23) 82.7 46.0 59.7 86.2 84.9
Adam-VPT [64] (Arxiv ’23) 70.8 34.8 53.9 84.0 81.1
Adam-FT [64] (Arxiv ’23) 65.7 48.5 56.1 86.5 84.4
Foster [52] (ECCV ’22) 87.3 - 5.1 86.9 -
iCARL [44] (CVPR ’17) 71.6 - 35.1 71.6 -
NCM [22] (NeurIPS-W ’22) 81.3 48.9 54.6 86.7 84.4
SLCA [62] (ICCV ’23) 86.8 - 54.2 82.1 -
RanPAC [30] (NeurIPS ’23) 89.6 ${}^{+}$ 67.3 ${}^{+}$ ${}^{+}$
RanDumb (Ours) 86.8 42.2 64.9 88.5 92.3
Table 9: RanDumb (Analysis). We thoroughly analyse the key components of RanDumb, and observe: (1) contributions of decorrelation and embedding, finding them interdependent and integral to performance; (2) consistent gains from augmentation across datasets; (3) marginal returns and saturation with higher embedding sizes, enabling computational tradeoffs; (4) increase in the optimal regularisation parameter for more complex datasets; (5) comparisons with alternate embeddings (or lack thereof) favoring RanDumb’s approach.
Method MNIST CIFAR10 CIFAR100 T-ImNet m-ImNet
(10/1) (10/1) (10/1) (200/1) (100/1)
Ablating Components of RanDumb
RanDumb 98.3 55.6 28.6 11.1 17.7
-Decorrelate 83.8 (-14.5) 30.0 (-25.6) 12.0 (-16.6) 4.7 (-6.4) 8.9 (-8.8)
-Embed 88.0 (-10.3) 41.6 (-14.0) 19.0 (-9.6) 8.0 (-3.1) 12.9 (-4.8)
-Both 82.1 (-16.2) 28.5 (-27.1) 10.4 (-18.2) 4.1 (-7.0) 7.28 (-10.4)
Effect of Adding Flip Augmentation
With - 55.6 28.6 11.1 17.7
Without 98.3 52.5 (-3.1) 26.9 (-1.7) 10.7 (-0.4) 16.6 (-1.1)
Variation with Ridge Parameter $\lambda$
$\lambda=10^{-6}$ 98.3 53.9 27.8 10.3 15.8
$\lambda=10^{-5}$ 55.6 28.6 11.1 15.9
$\lambda=10^{-4}$ 96.6 52.6 26.1 11.6 17.7
Variation Across Embedding Projections
No-Embed 88.0 41.6 19.0 8.0 12.9
RP+ReLU 95.2 48.8 23.1 9.7 15.7
RanDumb (Ours) 98.3 (+3.1) 55.6 (+6.8) 28.6 (+5.5) 11.1 (+1.4) 17.7 (+2.0)
3.2 RanDumb: Analysis
We extensively analyse RanDumb for the traditional OCL benchmarks in this subsection. First, we ablate the two components (random embedder and decorrelation) in RanDumb to ensure both components are necessary. Then we study effect of varying the embedding dimensionality, shrinkage and removing flipping augmentation and compare with alternate random embeddings proposed in literature [30] and across architectures [34].
Ablating Components of RanDumb. We ablate the contribution of only using Random Fourier features for embedding and decorrelation to the overall performance of RanDumb in Table 9 (top). Ablating the decorrelation and relying solely on random Fourier features, colloquially dubbed Kernel-NCM, has performance drops ranging from 6-25% across the datasets. Replacing random Fourier features with raw features, colloquially the SLDA baseline, leads to pronounced drop in performance ranging from 3-14% across the datasets. Moreover, ablating both components, resulting in the base nearest class mean classifier exhibits the poorest performance with an average reduction of 17%. The poor performance when ablating both decorrelation and embedding highlights their interdependence.
Figure 3: Impact of Embedding Dimensions. We show the variation of accuracy of RanDumb with embedding dimensionality across datasets. We see decreasing marginal returns with increasing dimensions, obtaining relatively minor improvements after 15K dimensions.
Impact of Embedding Dimensions. We vary the dimensions of the random Fourier features ranging from compressing 3K input dimensions to 1K to projecting it to 25K dimensions and evaluate its impact on performance in Figure 3. Surprisingly, the random projection to a 3x compressed 1K dimensional space allows for significant performance improvement over not using embedding, given in Table 9 (top). Furthermore, increasing the dimension from 1K to 25K results in improvements of 3.6%, 10.4%, 7.0%, and 2.5% on MNIST, CIFAR10, CIFAR100, and TinyImageNet respectively. Increasing the embedding sizes beyond 15K, however, only results in modest improvements of 0.1%, 1.4%, 1.1% and 0.2% on the same datasets, indicating 15K dimensions would be a good point for a performance-computational cost tradeoff.
Impact of Flip Augmentation. We evaluate the impact of adding the flip augmentation on the performance of RanDumb in Table 9 (middle). Note that MNIST was not augmented. Augmentation provided large gains of 3.1% on CIFAR10, 1.7% on CIFAR100, and 0.4% on TinyImageNet. We did not try further augmenting the data with RandomCrop transform as done with standard augmentations.
Impact of Varying Ridge Parameter. All prior experiments use a ridge parameter ($\lambda$) that increases with dataset complexity: $\lambda=10^{-6}$ for MNIST, $10^{-5}$ for CIFAR10 and CIFAR100, and $10^{-4}$ for TinyImageNet and miniImageNet. Table 9 (middle) shows the effect of varying $\lambda$ on RanDumb’s performance. With a smaller $\lambda=10^{-6}$, CIFAR10, CIFAR100, TinyImageNet and miniImageNet all exhibit minor drops of 0.1%-1.7%, 0.8%, 0.8%. Increasing shrinkage to a $\lambda=10^{-4}$ reduces CIFAR10 and CIFAR100 performance more substantially by 3% and 2.5% versus their optimal $\lambda=10^{-5}$. On the other hand, this larger $\lambda$ leads to improvements of 0.5% and 1.8% on TinyImageNet and miniImageNet. This aligns with the trend that datasets with greater complexity benefit from more regularisation, with the optimal $\lambda$ surfacing as a result of balancing under- and over-regularisation effects.
Comparison with RanPAC. We compared our random Fourier features with projection based on a random weight matrix and ReLU [30] (RP+ReLU) in Table 9 (middle) with their best embedding size. Our method performs significantly better on each dataset, averaging a gain of 3.4%. We believe this is attributable to using a theoretical grounded random projection compared to RanPAC [30], which is specifically designed for the setting using pretrained transformers.
Comparisons across Architectures. In table 7 (right), we compare whether using random Fourier features as embeddings outperforms models across various architectures for continual representation learning. We use experience replay (ER) baseline in the task-incremental CIFAR100 setup (for details, see Mirzadeh et al. [34] as it differs significantly from earlier setups). Our comparison involved across various architectures. The findings revealed that RanDumb surpassed the performance of nearly all considered architectures, and achieved close to 94% of the joint multi-task performance. This suggests that RanDumb outperforms continual representation learning across a wide range of architectures.
Conclusion. Overall, both random embedding and decorrelation are critical components in the performance of RanDumb. Using random fourier features is substantially better than RanPAC. Lastly, one can substantially reduce the embedding dimension without a large drop in performance for large gains in computational cost and RAM usage, additional augmentation may further significantly help performance and optimal shrinkage parameter increases with dataset complexity. RanDumb outperforms continual representation learning across a wide range of architectures.
4 Discussion and Concluding Remarks
Our investigation reveals a surprising result, empirically demonstrating that currently the representation gain is negative, i.e. $RepGain(\mathcal{D})<0$ — simply using random embedding (RanDumb) consistently outperforms learned representations from methods specifically designed for continual training. Furthermore, using fixed pretrained networks consistently outperforms continually finetuned method. Using random/pretrained features also recovers 70-90% of the gap to joint learning, leaving limited room for improvement in representation learning techniques. Our investigation is both surprising and alarming as it questions our understanding of how to effectively design and train models that require efficient continual representation learning, and necessitates a re-investigation of the widely explored problem formulation itself. We believe adoption of computationally bounded scenarios without memory constraints [41, 40, 17] could be a promising way forward.
Acknowledgements
AP is funded by Meta AI Grant No. DFR05540. PT thanks the Royal Academy of Engineering. PT and PD thank FiveAI for their support. This work is supported in part by a UKRI grant: Turing AI Fellowship EP/W002981/1 and an EPSRC/MURI grant: EP/N019474/1. The authors would like to thank Arvindh Arun, Kalyan Ramakrishnan and Shashwat Goel for helpful feedback.